OpenStudy (anonymous):

HELP MEDAL AND FAN HELP ASAP @ RANGA Identify whether the given value of the variable makes the equation true. –5r – 6 = –26; r = 4 A. false B. true

OpenStudy (anonymous):

HELP SOME ONE

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

HELP PLEASE

hero (hero):

Replace r with 4 in the formula, then simplify: -5(4) - 6 = ?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

that equals -26

hero (hero):

So it is true or false?

hero (hero):

False? Why do you believe it is false? Do you remember what the goal was here?

OpenStudy (anonymous):

sorry i came late computer laging

OpenStudy (anonymous):

Oh true sorry that is what i meant

OpenStudy (anonymous):

@Hero it is true

hero (hero):

Also, do not abuse the ping feature. Just ping one person. If they don't respond, then message the other users a link to your question.

hero (hero):

Yes it is true because -5(4) - 6 = -26 -20 - 6 = -26 -26 = -26 TRUE

OpenStudy (anonymous):

Oh i did not know that sorry

OpenStudy (anonymous):

@hero thanks

OpenStudy (anonymous):

@Reaper534 can you help with the next one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok i will tag you ok